what is the value of p in 6p-7=2p+7 and how did you get that answer?
Answer by jai_kos(139)

Given 6p - 7= 2p + 7
Group all the p terms on one side, we get
Whenever you bring one term to the other side, we change the sign.
6p - 2p = 7 + 7
4p = 14
Now divide by 4 on both sides, we get
p = 14 /4
p = 7 / 2
Therefore the value for p = (7 / 2)